OpenStudy (anonymous):

Need some geometry help please. Could someone help me with this problem Use the Distance Formula and the x-axis of the coordinate plane. Show why the distance between two points on a number line (the x-axis) is | a – b |, where a and b are the x-coordinates of the points. Thanks

OpenStudy (solomonzelman):

\(\Large\color{black}{ \displaystyle {\rm D} =\sqrt{ (\color{blue}{{\rm y}_1}-\color{red}{{\rm y}_2})^2+(\color{green}{{\rm x}_1}-\color{darkgoldenrod}{{\rm x}_2})^2 }}\) where \(\Large\color{black}{ \displaystyle {\rm D} }\) is the distance. \(\Large\color{black}{ \displaystyle (\color{green}{{\rm x}_1}~,~~\color{blue}{{\rm y}_1}) }\) and \(\Large\color{black}{ \displaystyle (\color{darkgoldenrod}{{\rm x}_2}~,~~\color{red}{{\rm y}_2}) }\) are your two points.

OpenStudy (solomonzelman):

your number line is the x, axis so your points are (a,0) and (b,0)

OpenStudy (solomonzelman):

And the definition of an absolute value for any function f(x) is \(\large\color{black}{ \displaystyle \sqrt{\left( f(x) \right)^2} }\)

OpenStudy (misty1212):

you are 15 and i am 22 the difference in our ages (the distance between 15 and 22) is \(|22-15|\) or \(|15-22|\)

OpenStudy (solomonzelman):

well we have to prove using distance formula.

OpenStudy (solomonzelman):

which I just (hopefully) showed how to do

OpenStudy (solomonzelman):

If there are questions, please ask. Don't leave or wait for snow/

OpenStudy (anonymous):

Sorry having site connection issues wont let me stay connected for more then a few seconds for some reason

OpenStudy (solomonzelman):

take your time:)

OpenStudy (anonymous):

So it would be starting like \[D = \sqrt{(0_{1}-0_{1})^{2}+ (a_{2}-b _{2})^{2} }\]

OpenStudy (solomonzelman):

well, not exactly, you don't need the sub symbols. it would just be \(\large\color{blue}{ \displaystyle {\rm D}=\sqrt{(0-0)^2+(a-b)^2} }\) where your points are \(\large\color{black}{ \displaystyle (a,~0) }\) and \(\large\color{black}{ \displaystyle (b,~0) }\).

OpenStudy (solomonzelman):

what can you simplify the expression inside the distance formula to?

OpenStudy (solomonzelman):

what is \(\large\color{blue}{ \displaystyle (0-0)^2 }\) equal to?

OpenStudy (anonymous):

It's 0

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

So, \(\large\color{blue}{ \displaystyle \sqrt{(0-0)^2+(a-b)^2} }\) is same as ?

OpenStudy (solomonzelman):

(same as square root of what ? )

OpenStudy (anonymous):

\[\sqrt{0 +ab } ?\] Sorry, really have not good with letters

OpenStudy (solomonzelman):

if you have struggle putting in the code in the equation editor you can always use the drawing tool:)

OpenStudy (anonymous):

No that's not it. Just not good with letters in math problems

OpenStudy (solomonzelman):

\({ \large\color{blue}{ \displaystyle {\rm D}=\sqrt{\color{red}{(0-0)^2}+(a-b)^2} } }\) \({ \large\color{blue}{ \displaystyle {\rm D}=\sqrt{\color{red}{0}+(a-b)^2} } }\) and you just remain with \({ \large\color{blue}{ \displaystyle {\rm D}=\sqrt{(a-b)^2} } }\)

OpenStudy (solomonzelman):

are we clear on this part (i posted just now) ?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

\(\large\color{blue}{{ \displaystyle \left|~\color{red}{f(x)}~\right|=\sqrt{(~\color{red}{f(x)}~)^2} } }\)

OpenStudy (solomonzelman):

this is the definition of the absolute value of any function (in fact this is how we differentiate it... later on in math) now, you have instead of \(\large\color{blue}{{ \displaystyle \color{red}{f(x)}} }\), the \(\large\color{blue}{{ \displaystyle \color{red}{a-b} } }\).

OpenStudy (solomonzelman):

so what is the expression in the root equivalent to?

OpenStudy (anonymous):

\[\left| a-b \right| = \sqrt{(a-b)^{2}}\]

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

and, wouldn't it be logical to say that \(\large\color{blue}{ \left| a-b\right|= \left| b-a\right| }\) ?

OpenStudy (solomonzelman):

oh, nvm you don't need that

OpenStudy (solomonzelman):

we got \(\large\color{blue}{ \left| a-b\right| }\) already

OpenStudy (solomonzelman):

so is everything clear so far, want a recap/review, have questions?

OpenStudy (anonymous):

I think I got it. So the answer would be \[D = \sqrt{(0-0)^{2} + (a - b)^{2}}\] \[D=\sqrt{0+(a-b)^{2}}\] \[D=\sqrt{(a-b)^{2}}\] \[\left| a-b \right|\]

OpenStudy (solomonzelman):

yes, those are the proper steps !!

OpenStudy (anonymous):

Thank you so much for your help

OpenStudy (solomonzelman):

yw